OpenStudy (anonymous):

HELP PLEASE!! Calculate probability of the independent event P(A or B) when P(a)=.3 and P(B)=.7

OpenStudy (anonymous):

You just have to multiply .3 by .7

OpenStudy (anonymous):

oh ok well what about conditional probability with the same variables?

OpenStudy (ybarrap):

$$ P(A\cup B)=P(A)+P(B)-P(A\cap B) $$ Since A and B are independent, \(P(A\cap B)=0\) So just SUM the independent probabilities.

OpenStudy (anonymous):

so just 1?

OpenStudy (ybarrap):

Specifically, this section - http://en.wikipedia.org/wiki/Inclusion%E2%80%93exclusion_principle#In_probability

OpenStudy (anonymous):

ok what about if the variables are P(a)=.5 P(b)=.4 and then P(a and b)=.1?????

OpenStudy (anonymous):

I'm supposed to calculate the probability

OpenStudy (ybarrap):

You've got everything you need, P(a), P(b) and P(a and b) -- see my formula above! Just plug in for P(a or b).

OpenStudy (ybarrap):

$$ P(A\cup B)=P(A)+P(B)-P(A\cap B)=0.5+0.4-0.1 $$ See how I did that?

OpenStudy (anonymous):

yes i just thought a and b were supposed to = P(A and B) but the answer is .8

OpenStudy (ybarrap):

Are you asking me?

OpenStudy (anonymous):

yes(:

OpenStudy (ybarrap):

What are you asking, not sure

OpenStudy (anonymous):

the answer is .8 right?

OpenStudy (ybarrap):

Yes for P(a or b).

OpenStudy (anonymous):

ok but theres nothing else i need to find? it just says calculate the probability

OpenStudy (ybarrap):

Tell me the whole question to be sure

OpenStudy (anonymous):

its 27

OpenStudy (ybarrap):

Oh, those are different problems that P(A or B) Ok

OpenStudy (ybarrap):

Do you know what P(A|B) means?

OpenStudy (anonymous):

i know its a conditional probability problem

OpenStudy (ybarrap):

$$ P(A|B)=\cfrac{PA\cap B}{P(B)}=\cfrac{0.1}{0.4} $$ You can do the next. BRB

OpenStudy (anonymous):

i got .25

OpenStudy (ybarrap):

Do #27 now

OpenStudy (anonymous):

umm isn't that what i did?

OpenStudy (anonymous):

oh okay for 27 i got 4

OpenStudy (anonymous):

if you can check the other ones i got .21 on 23 and on 25 i got 1 but if you can't then thank you for helping!

OpenStudy (ybarrap):

Probabilities can not be greater than 1 so #27 is not 4

OpenStudy (ybarrap):

For #27 what do you have for P(B|A)=P(B and A)/P(A) ?

OpenStudy (ybarrap):

We already did #23

OpenStudy (anonymous):

oh well you had .1/.4 so in 27 i just did .4/.1

OpenStudy (ybarrap):

No. The "|" does NOT mean division. It means "given" P(A|B) means Probability of event A given event B.

OpenStudy (anonymous):

you did .1/.4

OpenStudy (ybarrap):

$$ P(B|A)=\cfrac{P(A\cap B)}{P(A)}\\ P(A|B)=\cfrac{P(A\cap B)}{P(B)} $$ Do you see the difference above between the two?

OpenStudy (ybarrap):

Yes and that one is correct

OpenStudy (anonymous):

i guess I'm not understanding this then

OpenStudy (ybarrap):

So we are dealing with probabilities of events. When we are given that an event occurs, we are asked what is the probability that the other will occur. If they are independent, then there is no effect. But if they are not independent, then conditional probability tells us that Probability of A given B is the Probability of A and B divided by the probability of B. Here is a visual 

OpenStudy (anonymous):

ok

OpenStudy (ybarrap):

That fraction is $$ P(A|B)=\cfrac{P(A\cap B)}{P(B)} $$

OpenStudy (anonymous):

that looks like alien writing to me

OpenStudy (ybarrap):

$$ P(A|B)=P(A \text{ given }B)=\cfrac{P(A\cap B)}{P(B)}=\cfrac{P(A\text{ and }B)}{P(B)} $$ Is this better?

OpenStudy (anonymous):

haha no I'm sorry... what does it look like with the values plugged in

OpenStudy (ybarrap):